Consider the titration of 100 ml of 1M NaF with 1M HCl. Calculate the initial pH.
Answer #1

F- partly hydrolysed by water

F-(aq) + H2O(l) < ------- > HF(aq) + OH-(aq)

Kb = [HF][OH-] / [F-]

Kb = Kw/Ka

where,

Kw = ionic product of water , 1.00 × 10-14

Ka = Ka of HF , 6.92 × 10-4

Kb = 1.00 ×10-14/ 6.92 ×10-4

Kb = 1.45 × 10-11

Initial concentration

[F-] = 1

[HF] = 0

[OH-] = 0

change in concentration

[F-] = -x

[HF] = +x

[OH-] = + x

Equilibrium concentration

[F-] = 1- x

[HF] = x

[OH-] = x

so,

x2/(1 - x) = 6.92 ×10-4

solving for x

x = 0.02596

[OH-] = 0.02596M

pOH = -log[OH-]

pOH = -log(0.02596)

pOH = 1.59

pH = 14 - pOH

pH = 14 - 1.59

pH = 12.41

Therfore

Initial pH = 12.41